A room is in the shape of a rectangle. The width of the room is 2 feet longer than two thirds the length. If the room is 33 feet long, what is the area?
Answer by CubeyThePenguin(3113)

width = 2/3 * (33) + 2 = 24 ft

area = (24)(33) = 792 ft^2